OpenStudy (anonymous):

I have a guestion Suppose that Jamie bought a bike in 2009 and sold it for a loss of 18% a year later. If she sold it for $116.64 what was theioriginal purchase price?

OpenStudy (amistre64):

116.64 = original price - original price(18%)

OpenStudy (amistre64):

116.64 = P - P(.18) 166.64 = P(1 - .18) P = 166.64 /(1-.18)

OpenStudy (amistre64):

203.21951219512195121951219512195

OpenStudy (anonymous):

price is 116.54

OpenStudy (amistre64):

then just put the right price into the equation...and work it :)

OpenStudy (anonymous):

ok sorry the price is 116.64 I can type today

OpenStudy (amistre64):

oh ...yeah, that 166.64 is just a typo in my head lol :)

OpenStudy (amistre64):

P = 116.64/(1-.18)

OpenStudy (anonymous):

so it the answer 203.22

OpenStudy (amistre64):

142.2439 is the correct answer with everything put in as intended :)

OpenStudy (anonymous):

that is what I got also thanks so much